OpenStudy (anonymous):

Solve the following system of equations by using the elimination method. x + y = 4 2x + 3y = 9

OpenStudy (anonymous):

(3,1) is wat i got but i dont know if it is right

OpenStudy (rsadhvika):

to knw if its right or not, plug them back and test

OpenStudy (rsadhvika):

(3, 1) x+y = 4 3+1 = 4 4 = 4 true

OpenStudy (rsadhvika):

2x+3y = 9 2(3) + 3(1) = 9 6+3 = 9 9=9 true

OpenStudy (rsadhvika):

so (3, 1) satisfies both equations. so wat u got is right ! good job !!